If npgreater than or equals≥5 and nqgreater than or equals≥​5, estimate Upper P left parenthesis more...

If

npgreater than or equals≥5

and

nqgreater than or equals≥​5,

estimate

Upper P left parenthesis more than 5 right parenthesisP(more than 5)

with

nequals=1111

and

pequals=0.40.4

by using the normal distribution as an approximation to the binomial​ distribution; if

npless than<5

or

nqless than<​5,

then state that the normal approximation is not suitable.

Select the correct choice below​ and, if​ necessary, fill in the answer box to complete your choice.

A.

Upper P left parenthesis more than 5 right parenthesisP(more than 5)equals=nothing

​(Round to four decimal places as​ needed.)

B.

The normal distribution cannot be used.
Answer #1

Given :-

n=11

p=0.4

q = 1-p = 1-0.4 = 0.6

np = 11 * 0.4 = 4.4 < 5

nq = 11 * 0.6 = 6.6 ≥5

First conditions are not satisfied. So, the normal approximation is not suitable.

##Answer :- option "B" is correct.

The normal distribution cannot be used.